OpenStudy (anonymous):

In a large bag of Skittles®, each of the 5 colors (red, orange, yellow, blue, green) occurs with the same probability. You reach in and select 1 candy. Find P(a color in the U.S. flag).

OpenStudy (anonymous):

how would i figure this out?

OpenStudy (anonymous):

P(a color in the U.S. flag) = P(red OR blue) = P(red) + P(blue) - P(Red AND Blue) P(Red and Blue) = 0 since you can not pick both at the same time P(Any color) = 1/5 P(red) = 1/5 P(Blue) = 1/5 P(Red OR Blue) = 1/5 + 1/5 = 2/5